Skysong Company uses a job costing system. Overhead is allocated based on 120% of direct labour cost. Last months transactioWhat is the amount of direct materials used for Job 850? Direct materials
Answer #1

a) Ending WIP = Beginning WIP+Direct material+Direct labor+Overhead-Ending WIP

= 46100+162500+113300+151700-446400

Ending WIP = 27200

b) Direct labor = 9600/1.2 = 8000

c) Direct material = Ending WIP-Overhead-Direct labor = 27200-9600-8000 = 9600